OpenStudy (anonymous):

which equation represents a line that is parallel to the line represented by the equation y=3/4x-1/2 a)y=2x-3/4 b)y=3/4x+1/2 c)y=-4/3x-1/2 d)y=-1/2+3/4

OpenStudy (anonymous):

y=mx+b If the m is the same, the lines are parallel.

OpenStudy (anonymous):

It is b) y=3/4x+1/2 which has the same slope 3/4 as that of the given line y=3/4x-1/2 and so is the line parallel to the line y=3/4x-1/2